OpenStudy (sleepyjess):

Find the vertex, focus, directrix, and focal width of the parabola. x = 4y^2

OpenStudy (sleepyjess):

@wio could you help on this one?

OpenStudy (alexandervonhumboldt2):

first isolate y

OpenStudy (sleepyjess):

Vertex is 0, 0 , but I am getting confused on how to get the focus

OpenStudy (alexandervonhumboldt2):

hmm yeah i dont know too

OpenStudy (alexandervonhumboldt2):

@wio

OpenStudy (sleepyjess):

or @satellite73 ?

OpenStudy (anonymous):

Well, go back to the equation: \[ (x-x_0)^2=4p(y-y_0) \]But we want to change it to: \[ (y-y_0)^2=4p(x-x_0) \]In this case.

OpenStudy (anonymous):

First:\[ y^2=4px \]And then \[ x = 4y^2\implies y^2=\frac x4 \]And then \[ 4px = \frac x4 \implies 4p=\frac 14\implies p=\frac 1{16} \]

OpenStudy (sleepyjess):

That helps so much! :) Thanks again

OpenStudy (mathmate):

We'll get the nomenclature clear  It can be seen from the diagram that for a horizontal axis parabola where vertex is at the origin, y^2=4px, or p=1/16 to give x=4y^2 as @wio had it. since p=1/16, so the directrix is at x=-1/16, and the focus is at (+1/16, 0). The focal width, is the distance between the two intersections of the line parallel to the directrix, but passing through the focus. In this case, this line is x=1/16, which gives y^2=4px=4x/16=1/64 so y=\(\pm 1/8\), and the focal width, or latus rectum is 2(1/8)=1/4.

OpenStudy (mathmate):